12 movies - 5 action, 5 romantic comedies, 2 thrillers 6 snacks - 1 popcorn, 3 chips, 2 cookies
4 drinks - 2 juice, 2 soda
How many different combinations can I make to select 2 movies, 2 snacks, and 1 drink, if the order of the movies does not matter?

Answer by greenestamps(13209)



Choose 2 of the 12 movies: 12 choose 2 = (12*11)/(2*1) = 66
Choose 2 of the 6 snacks: 6 choose 2 = (6*5)/(2*1) = 15
Choose 1 of the 4 drinks: 4 choose 1 = 4

Choose 2 of the 12 movies AND 2 of the 6 snacks AND 1 of the 4 drinks = 66*15*4 = 3960